# Robust estimates of theoretical uncertainties at fixed-order in perturbation theory

### v1

@@ I Introduction
Figure 1: Comparison between MHOU estimates from TNP and scale variation for differential distributions in $t\bar{t}$ production, with QCD corrections to the production and decay processes treated separately and combined in the narrow width approximation. From left to right, we show the lepton pair invariant mass $m(\ell\bar{\ell})$, the transverse momentum of the hardest $b$-jet $p_{T}(b_{1})$ and the ratio of the lepton transverse momenta $p_{T}(\ell_{1})/p_{T}(\ell_{2})$. We take $\mu=H_{T}/4$ as our central scale choice and use a Bernstein parameterisation. The solid lines represent the LO (green), NLO (blue) and NNLO (red) central predictions and the corresponding bands of the estimated MHOU. The vertical lines indicate statistical uncertainties. We do not associate a TNP uncertainty with the LO prediction – see text for details.
@@ II Implementation
where $f_{k}$ is a polynomial of degree $k$ in a (suitably rescaled) kinematic variable $x$ and the $\theta_{i}$ are the TNPs. We have experimented with different choices of polynomial basis, and present results using Bernstein polynomials:
@@
Figure 3: As in fig. 1, but for off-shell, fully decayed $WW$ production. We take $\mu=H_{T}/2$ as our central scale choice and use a Bernstein parameterisation. From left to right we show the pair invariant mass $m(W^{+}W^{-})$, the transverse momentum of the muon $p_{T}(\mu^{-})$ and the rapidity of one boson $y(W^{-})$.
@@
We note that strict adherence to the philosophy espoused in Ref. [5] would dictate that our model be independent of the size (if not the shape) of lower-order corrections. The motivation for this is to avoid significant underestimation of the MHOU when the lower-order terms happen by chance to be smaller than might be naïvely expected. While this is certainly a desirable property in principle, in practice we have found it difficult to achieve, even after factoring in known structures (flavour decomposition of partonic channels, etc.). Since our aim here is to provide a prescription which can be applied generally and to a range of processes, we have conceded on this particular tenet [^2] In this sense, our approach is close to that adopted in Ref. [9], where the authors also use the size of lower-order terms in their modelling.. We note that in so doing, we do not observe any instances where we believe the MHOU has been significantly underestimated (in contrast to results obtained using normal scale variations). While we do not exclude the possibility that it is possible to achieve a parameterisation which does not utilise this size information, its form is likely to be specific to a given interest case, given that it depends on a complex interplay of factors (e.g. partonic channels, the definition of fiducial regions, isolation criteria).
@@ III Results
We obtain our NLO and NNLO cross sections from published results obtained within the STRIPPER framework [10, 11, 12, 13], a full list of included processes can be found in Tab. 1. We study both NLO uncertainties (where the NNLO corrections are known) and NNLO uncertainties (for cases where the N3LO corrections are at present unknown). We explicitly present here results for three processes, namely $pp\to t\bar{t}\to b\bar{b}\ell\bar{\ell}$, $pp\to\gamma jj$ and $pp\to W^{+}W^{-}\to\ell\bar{\ell}\nu\bar{\nu}$, all defined in appropriate fiducial regions [^3] More processes are available using a Jupyter notebook interface to HighTEA at Ref. [14].. Each case has a distinct behaviour with respect to scale variations. The first and second processes feature coloured final states and large higher-order corrections, where typically the leading-order scale band does not contain the NLO result. The third example is pathological with respect to scale variations, and the NNLO uncertainty band lies completely outside that of the NLO. This allows us to test our TNP procedure in different scenarios. For the original NNLO computations of each process, see Refs. [15, 16, 17, 18].
@@
For each process and distribution we consider, we compute the central curve using a fixed scale choice. We provide uncertainty estimates computed in two different ways. In the first method, we generate random values of the $\theta_{i}$ uniformly in the interval $[-1,1]$ and take a band of the (bin-wise) maximum and minimum deviations thus obtained.  [^4] We have verified that the size of the band is independent of the number of samples above $N=100$. We note that this latter part of the procedure is non-ideal if one wishes to propagate the uncertainties correctly – each set of the $\theta_{i}$ essentially represents a correlation model, and by taking a band one loses information. For our purposes however, it is serves as a useful tool to demonstrate the comparison with a scale variation method. In the second method, we sum in quadrature variations by ±$\pm 1$ of each nuisance parameter separately. This is appropriate if the parameters are independently and normally distributed, a matter which we shall address in the following. We show results using both Bernstein and Chebyshev parameterisations depending on the process – we have verified that the results are very similar, no matter which is used.
@@
We begin by discussing results for the $t\bar{t}$ process with a central scale choice of $H_{T}/4$. We show the invariant mass of the lepton pair $m(\ell\bar{\ell})$, the transverse momentum of the hardest $b$-jet $p_{T}(b_{1})$ and the ratio of the lepton transverse momenta $p_{T}(\ell_{1})/p_{T}(\ell_{2})$. From the lower scale variation plots we see that the LO uncertainty is generally underestimated but that beyond this order, scale variations appear to estimate the MHOU well for all spectra. The TNP approaches reproduce this behaviour, with marginally smaller estimates of the NLO uncertainty. The band approach seems in general to give slightly more conservative estimates than the quadrature method.
Next, we examine the case of $\gamma jj$ production at a central scale of $H_{T}$. Examining first the lower plots, we note that, as in the $t\bar{t}$ case, the LO scale variation band does not contain the nominal prediction at NLO for any of the four distributions presented. Thereafter, scale variations again seem to estimate the MHOU well, as do both TNP approaches. We stress here that the effect of the TNP variations is not only an increase in the size of the bands, but also a modification of their shape. This is most clearly visible at NNLO in the $p_{T}^{j}$ and $|\Delta y^{\gamma-j}|$ distributions, where the scale variations give an asymmetric uncertainty while the TNP variation is more even.
Finally, we turn to off-shell $WW$ production. From the lower scale variation plots we see that the uncertainty at each order is completely underestimated. This is a result of so-called ‘giant’ $K$-factors, which arise from NLO corrections in which a vector boson is emitted off a final-state parton [19, 20, 21]. The size of these corrections is exacerbated by fiducial cuts. In the case of the invariant mass distribution, the LO band is in fact smaller that the NLO which is of a similar size to the NNLO scale variation. In none of the distributions do the scale bands even overlap. In contrast, while the process is clearly challenging even for the TNP approach, a much more sensible behaviour is observed. The TNP estimates decrease in size as one moves from NLO to NNLO, and the NLO error in particular seems to be much better estimated. The TNP band approach in particular shows a much improved behaviour – while the NNLO curve is not perfectly contained within the NLO band, it easily lies within $2\sigma$.
@@
Figure 5: Distribution of the values of the nuisance parameters $\theta_{i}$ in a Bernstein parameterisation at NNLO. For details of the data points included, see text. The mean and standard deviation of a fit Gaussian distribution are indicated on the plot. Binning is purely illustrative.
@@
As an a posteriori check on our procedure, we have performed fits of our NLO uncertainty estimate in cases where the NNLO distribution is known. An example of such a fit for the process $pp\to ZZ^{*}\to e^{+}e^{-}\mu^{+}\mu^{-}$ is shown in Fig. 4. This confirms that our parameterisation is able to capture the true behaviour of the NNLO calculation with values of $\theta_{i}\sim\mathcal{O}(1)$.
@@
At this point, a comment is in order. It is true that, without knowledge of the functional form of the higher-order corrections (which one possesses, for example, in a resummed calculation), it is not possible to capture the correct bin-to-bin correlations of the uncertainties in a given differential distribution. This is only possible to the extent that the parameterisation in eq. 2 provides a good approximation of the true higher-order behaviour. Given that this ‘true’ behaviour is in practice always accompanied by a finite Monte Carlo error, the quality of the fits in Fig. 4 demonstrate that this is indeed the case: that is, the parameterisation describes correlations as well as the integration error allows. It is of course possible that by increasing statistics, one might find that a polynomial with $k=2$ no longer describes the functional form well, and in fact underfits. In this case, one could always extend the parameterisation by choosing $k>2$.
@@
We have repeated the fit for a range of processes and distributions, which are listed in the Tab. 1. In selecting the kinematic distributions to be included, we have endeavoured to make choices which are quasi-independent – for example, choosing a rapidity distribution, an invariant mass and a transverse momentum rather than two invariant masses. In this way, we have tried to avoid double-counting information arising from strongly correlated distributions. This also restricts the amount of information we are able to include from each process – for $2\to 2$ processes we limit ourselves to three distributions, while more can be included in the $2\to 3$ cases due to the greater kinematic freedom they afford. In total, we are able to include 30 data points.
Figure 6: As in Fig. 5 but for Chebyshev polynomials.
@@
Results for the fit values of the nuisance parameters are shown in figs. 5 and  6, for both Bernstein and Chebyshev parameterisations. We observe that the resulting distributions appear to be Gaussian, with values of the mean and standard deviation of order 1. The test statistics obtained using the Shapiro-Wilk and Anderson-Darling methods show no significant deviation from normality at the $p=0.05$ level. This is a strong justification for our procedure, at least at NLO – a similar validation at NNLO is unfortunately dependent on a greater availability of fully differential calculations at N3LO.
@@ IV Conclusions
It would be interesting to apply our method to obtain uncertainties for differential distributions at N3LO, and to compare these to the usual scale variation estimates. At present, the only fully-differential calculations at N3LO are for Higgs boson production in gluon fusion [34] and Drell-Yan production [35]. Our method could also be used in the context of electroweak corrections. These present a problem to scale variation procedures due to the weak dependence on $\mu$ of the electroweak coupling. A nuisance parameter approach, on the other hand, would not suffer from this problem and could thus improve uncertainty estimates in a consistent fashion.
@@ Acknowledgements
We are grateful to Frank Tackmann and Thomas Cridge for discussions on this topic and for comments on the manuscript. We thank Tom Schellenberger for preparing the numerical results for Higgs production. This work was initiated at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452 and by a grant from the Simons Foundation (1161654, Troyer). MAL is supported by the UKRI guarantee scheme for the Marie Skłodowska-Curie postdoctoral fellowship, grant ref. EP/X021416/1. RP wishes to thank the CERN Theoretical Physics Department for hospitality while part of this work was carried out.

### v2

@@ I Introduction
Figure 1: Comparison between MHOU estimates from TNP and scale variation for differential distributions in $t\bar{t}$ production, with QCD corrections to the production and decay processes treated separately and combined in the narrow width approximation. From left to right, we show the lepton pair invariant mass $m(\ell\bar{\ell})$, the transverse momentum of the hardest $b$-jet $p_{T}(b_{1})$ and the ratio of the lepton transverse momenta $p_{T}(\ell_{1})/p_{T}(\ell_{2})$. We take $\mu=H_{T}/4$ as our central scale choice and use a Bernstein parameterisation. The solid lines represent the LO (green), NLO (blue) and NNLO (red) central predictions and the corresponding bands of the estimated MHOU from scale variations, the dashed bands those from TNPs. The vertical lines indicate statistical uncertainties. We do not associate a TNP uncertainty with the LO prediction – see text for details.
@@ II Implementation
where $f_{k}$ is a polynomial of degree $k$ in a (suitably rescaled) kinematic variable $x$ and the $\theta_{i}$ are the TNPs. [^1] In the notation of Ref. [5], this is then an Nm+1LO model for a fixed value of the scale $\mu$. We have experimented with different choices of polynomial basis, and present results using Bernstein polynomials:
@@
Figure 3: As in fig. 1, but for off-shell, fully decayed WWWW production. We take $\mu=H_{T}/2$ as our central scale choice and use a Bernstein parameterisation. From left to right we show the pair invariant mass $m(W^{+}W^{-})$, the transverse momentum of the muon $p_{T}(\mu^{-})$ and the rapidity of one boson $y(W^{-})$.
Figure 4: Comparison between scale and TNP uncertainties for charged-current ($W^{+}$) and neutral-current Drell-Yan ($Z/\gamma$) through next-to-next-to-next-to-leading order. The coloured bands indicate the uncertainties from 7-point scale variations, and dotted lines the quadrature combination of the TNP uncertainties at a 95% CL.
@@
We note that strict adherence to the philosophy espoused in Ref. [5] would dictate that our model be independent of the size (if not the shape) of lower-order corrections. The motivation for this is to avoid significant underestimation of the MHOU when the lower-order terms happen by chance to be smaller than might be naïvely expected. While this is certainly a desirable property in principle, in practice we have found it difficult to achieve, even after factoring in known structures (flavour decomposition of partonic channels, etc.). Since our aim here is to provide a prescription which can be applied generally and to a range of processes, we have conceded on this particular tenet [^3] The closest analogue in Ref. [5] would be strategy 3 of section 4.3. In that case, however, the author is explicit about the normalisation required to make the estimate independent of the size of lower order terms. In using this information, our approach is instead closer to that adopted in Ref. [9].. We note that in so doing, we do not observe any instances where we believe the MHOU has been significantly underestimated (in contrast to results obtained using normal scale variations). While we do not exclude the possibility of achieving a parameterisation which does not utilise this size information, its form is likely to be specific to a given interest case, given that it depends on a complex interplay of factors (e.g. partonic channels, the definition of fiducial regions, isolation criteria).
@@ III Results
We obtain our (N)NLO cross sections from published results obtained within the STRIPPER framework [10, 11, 12, 13]: a full list of included processes can be found in Tab. 1. We study both NLO uncertainties (where NNLO corrections are known) and NNLO uncertainties (where N3LO corrections are at present unknown). We explicitly present results for three processes, namely $p\to t\bar{t}\to b\bar{b}\ell\bar{\ell}, pp$$p\to\gamma jj$ and $pp\to W^{+}W^{-}\to\ell\bar{\ell}\nu\bar{\nu}$, all defined in appropriate fiducial regions. More processes are available using a Jupyter notebook interface to HighTEA at Ref. [14]. Each case has a distinct behaviour with respect to scale variations. The first and second processes feature coloured final states and large higher-order corrections, where typically the leading-order scale band does not contain the NLO result. The third example is pathological with respect to scale variations, and the NNLO uncertainty band lies completely outside that of the NLO. This allows us to test our procedure in different scenarios. For the original NNLO computations of each process, see Refs. [15, 16, 17, 18].
@@
For each process and distribution we consider, we compute the central curve using a fixed scale choice. We compute uncertainty estimates by summing in quadrature variations by $\pm 1$ of each nuisance parameter separately. This is appropriate if the parameters are independently and normally distributed, a matter which we shall address in the following. We show results using both Bernstein and Chebyshev parameterisations depending on the process – we have verified that the results are very similar, no matter which is used.
@@
We begin by discussing results for the $t\bar{t}$ process with a central scale choice of $H_{T}/4$. In Fig. 1 we show the invariant mass of the lepton pair $m(\ell\bar{\ell})$, the transverse momentum of the hardest $b$-jet $p_{T}(b_{1})$ and the ratio of the lepton transverse momenta $p_{T}(\ell_{1})/p_{T}(\ell_{2})$. From the scale variation plots we see that the LO uncertainty is generally underestimated but that beyond this order, scale variations appear to estimate the MHOU well. The TNP approach reproduce this behaviour, with marginally smaller estimates of the NLO uncertainty.
@@
Figure 6: Distribution of the values of the nuisance parameters $\theta_{i}$ in Bernstein and Chebyshev parameterisations at NNLO. For details of the data points included, see text. The mean and standard deviation of a fit Gaussian distribution are indicated on the plot, as is the Anderson-Darling test statistic (critical value 0.712 for $p=0.05$). Binning is purely illustrative.
Next, we examine in Fig. 2 the case of $\gamma jj$ production at a central scale of $H_{T}$. Examining first the lower plots, we note that, as in the $t\bar{t}$ case, the LO scale band does not contain the nominal prediction at NLO for any of the four distributions presented. Thereafter, scale variations again seem to estimate the MHOU well, as does the TNP approach. We stress here that the effect of the TNP variations is not only an increase in the size of the uncertainties, but also a modification of their shape. This is most clearly visible at NNLO in the $p_{T}^{j}$ and $|\Delta y^{\gamma-j}|$ distributions, where the scale variations give an asymmetric uncertainty while the TNP variation is more even.
In Fig. 3 we turn to off-shell $WW$ production, a process notorious for the poor behaviour of traditional methods in estimating MHOU terms. In this preliminary study, we do not include the effect of the loop-induced $gg$ channel at NNLO. From the lower scale variation plots we see that the uncertainty at each order is completely underestimated. This is a result of so-called ‘giant’ $K$-factors, which arise from NLO corrections in which a vector boson is emitted off a final-state parton [19, 20, 21]. The size of these corrections is exacerbated by fiducial cuts. In the case of the invariant mass distribution, the LO band is in fact smaller that the NLO which is of a similar size to the NNLO scale variation. In none of the distributions do the scale bands even overlap. In contrast, while the process is clearly challenging even for the TNP approach, a much more sensible behaviour is observed. The TNP estimates decrease in size as one moves from NLO to NNLO, and the NLO error in particular seems to be much better estimated.
Finally, we discuss the case of neutral and charged current Drell-Yan production. This is a rare case where fully-differential corrections in QCD are currently known through N3LO [22, 23] – indeed, in Refs. [24, 25] the first predictions at this order for the invariant mass distribution of the dilepton pair were presented. It was observed that, beginning at N3LO, the convergence of the perturbative series as measured by scale variation begins to deteriorate. While this is partly due to the effect of using NNLO PDFs (as opposed to aN3LO [26]), it is nevertheless interesting to examine whether TNPs provide an uncertainty estimate which suggests greater consistency between perturbative orders. In Fig. 4 we show predictions for the invariant mass distribution at various orders in perturbation theory, for both charged and neutral current Drell-Yan production, and compare scale variations with (Chebyshev) TNP uncertainty estimates at the 95% confidence level.  [^4] This corresponds to varying the TNPs by 2, rather than 1 as done elsewhere. We wish to stress that, given the statistical interpretation of the TNPs which we will demonstrate in the next section, this genuinely corresponds to a $2\sigma$ uncertainty – it is thus not analogous to a variation of $\mu$ by a factor of 4. Results have been obtained using the public code n3loxs [27]. A clear improvement of the description of the NNLO uncertainty in the TNP case with respect to the scale variation case is visible, particularly for $Q>100~\mathrm{GeV}$ where the uncertainties now enclose the N3LO result. We are thus able to employ our method, without modification and in a simple and straightforward manner, to resolve an important discrepancy at the highest available order in perturbation theory.
@@
As an a posteriori check on our procedure, we have performed fits of our NLO uncertainty estimate in cases where the NNLO distribution is known. An example of such a fit for the process $pp\to ZZ^{*}\to e^{+}e^{-}\mu^{+}\mu^{-}$ is shown in Fig. 5. This confirms that our parameterisation is able to fit the NNLO data with values of $\theta_{i}\sim\mathcal{O}(1)$.
@@
At this point, a comment is in order. It is true that, without knowledge of the functional form of the higher-order corrections (which one possesses, for example, in a resummed calculation), it is not possible to capture the correct bin-to-bin correlations of the uncertainties in a given differential distribution. This is only possible to the extent that the parameterisation in eq. 2 provides a good approximation of the true higher-order behaviour. Given that this ‘true’ behaviour is in practice always accompanied by a finite Monte Carlo error, the quality of the fits in Fig. 5 demonstrate that this is indeed the case: that is, the parameterisation describes correlations as well as the integration error allows. It is of course possible that by increasing statistics, one might find that a polynomial with $k=2$ no longer describes the functional form well, and in fact underfits. In this case, one could always extend the parameterisation by choosing k>2k>2555We have studied the effect of truncation uncertainties using a toy model and simulated noise. Our expectation is that the induced truncation error is currently small compared to the nominal uncertainty estimate. Alternatively, it may be the case that by including too many TNPs one overfits. The treatment here depends on the exact parameterisation chosen. In the Chebyshev case, the orthogonality of the polynomial basis results in higher order coefficients naturally becoming smaller. In the Bernstein case, correlations between the TNPs are introduced which effectively set the coefficients of higher order monomials to zero. Care must therefore be taken when combining uncertainties or using them in a fit.
@@
We have repeated the fit for a range of processes and distributions, which are listed in Tab. 1. In selecting the kinematic distributions to be included, we have endeavoured to make choices which are quasi-independent – for example, choosing a rapidity distribution, an invariant mass and a transverse momentum rather than two invariant masses. In this way, we have tried to avoid double-counting information arising from strongly correlated distributions. This also restricts the amount of information we are able to include from each process – for $2\to 2$ processes we limit ourselves to three distributions, while more can be included in the $2\to 3$ cases due to the greater kinematic freedom they afford. In total, we are able to include 30 data points.
@@
Results for the fit values of the nuisance parameters are shown in figs. 6, for both Bernstein and Chebyshev parameterisations. We observe that the resulting distributions appear to be Gaussian, with values of the mean and standard deviation of order 1. Test statistics obtained using the Shapiro-Wilk and Anderson-Darling methods show no significant deviation from normality at the $p=0.05$ level. This is a strong justification for our procedure, at least at NLO – a similar validation at NNLO is unfortunately dependent on a greater availability of fully differential calculations at N3LO.
@@ IV Conclusions
Our method could also be used in the context of electroweak corrections. These present a problem to scale variation procedures due to the weak dependence on $\mu$ of the electroweak coupling. A nuisance parameter approach, on the other hand, would not suffer from this problem and could thus improve uncertainty estimates in a consistent fashion.
In practical applications, one is often interested in capturing not only the correlations in uncertainties between bins of a given kinematic distribution as is done here, but also between different distributions for a given process. Combining information in this way can be used to extract Standard Model parameters precisely, see e.g. ref. [40, 41]. Such a combination would require a multi-dimensional TNP treatment, which parameterises the joint dependence in multiple variables $x_{i}$. A simple attempt at this could for example use products of single-differential parameterisations. We leave a more detailed investigation of this topic to future work. Achieving the correct correlation model for differential distributions between processes would instead involve a significant departure from the approach we have adopted here and will require substantially new ideas.
@@ Acknowledgements
We are grateful to Frank Tackmann and Thomas Cridge for discussions on this topic and for comments on the manuscript. We thank Tom Schellenberger for preparing the numerical results for Higgs production. This work was initiated at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452 and by a grant from the Simons Foundation (1161654, Troyer). MAL is supported by the UKRI guarantee scheme for the Marie Skłodowska-Curie postdoctoral fellowship, grant ref. EP/X021416/1. RP thanks the CERN Theoretical Physics Department for hospitality while part of this work was carried out.